# The Geometry of Concepts: Sparse Autoencoder Feature Structure

### v1

@@ 5.1 Shape analysis
Figure 6 (left) quantifies this by showing the eigenvalues of the point cloud’s covariance matrix in decreasing order. revealing that they are not constant, but appear to fall off according to a power law. To test whether this surprising power law is significant, the figure compares it with the corresponding eigenvalue spectrum for a point cloud drawn from an isotropic Gaussian distribution, which is seen to be much flatter and consistent with the analytic prediction: the covariance matrix of $N$ random vectors from a multivariate Gaussian distribution follow a Wishart Distribution, which is well studied in random matrix theory. Since the abrupt dropoff seen for the smallest eigenvalues is caused by limited data and vanishes in the limit $N\to\infty$, we dimensionally reduce the point clound to its 100 largest principal componends for all subsequent analysis in this section. In other words, the point cloud has the shape of a “fractal cucumber”, whose width in successive dimensions falls off like a power law. We find such power law scaling is significantly less prominent for activations than for SAE features; it will be interesting for further work to investigate its origins.
Figure 7: Eigenvalues of the point cloud are seen to decay as an an approximate power law (left), whose slope depends on layer (right) and is strongly inconsistent with sampling from an isotropic Gaussian distribution. The effective point cloud volume is also seen to be layer-dependent (right).
Figure 6 (right) shows how the slope of the aforementioned power law depends on LLM layer, computed via linear regression against the 100 largest eigenvalues. We see a clear pattern where middle layers have the steepest power law slopes: (Layer 12 has slope -0.47, while early and late layers (e.g., Layers 0 and 24) have shallower slopes (-0.24 and -0.25), respectively. This may hint that middle layers act as a bottleneck, compressing information into fewer principal components, perhaps optimizing for more efficient representation of high-level abstractions. Figure 7 (right) also shows how effective cloud volume (the determinant of the covariance matrix) depends on layer, on a logarithmic scale.
@@ 5.2 Clustering analysis
Figure 8: Estimated clustering entropy across layers with 95% confidence intervals. Middle layers exhibit reduced clustering entropy, while earlier and later layers show higher entropy, indicating distributed and concentrated feature representation in early and late layers, respectively.
Clustering of galaxies or microscopic particles is often quantified in terms of a power spectrum or correlation function. This is complicated for our very high-dimensional data, since the underlying density of varies with radius and, for a high-dimensional Gaussian distribution, is strongly concentrated around a relatively thin spherical shell. For this reason, we instand quantify clustering by estimating the entropy of the distribution that the point cloud is assumed to be sampled from. We estimate the entropy $H$ from our SAE feature point cloud using on the $k$-th nearest neighbor (k-NN) method Dasarathy (1991); Kozachenko & Leonenko (1987), computed as follows,
@@
where $r_{i}$ is the distance to the $k$-th nearest neighbor for point $i$, and $d$ is the dimensionality of the point cloud; $n$ is the number of points; the constant $\Psi$ is the digamma term from the k-NN estimation. As a baseline, the Gaussian entropy represents the maximum possible entropy for a given covariance matrix. For a Gaussian distribution with the same covariance matrix, the entropy computed as:
@@
where $\lambda_{i}$ are the eigenvalues of the covariance matrix. We define the clustering entropy (often referred to as “negentropy” in physics as $H_{gauss}-H$, i.e., how much lower the entropy is than its maximum allowed value. Figure 8 shows the estimated clustering entropy across different layers. We see that the SAE point cloud is strongly clustered, particulary in the middle layers. In future work, it will be interesting to examine whether these variations depend mainly on the prominence of crystals or lobes in different layers, or have an altogether different origin.
6 Conclusion
In this paper, we have found that the concept universe of SAE point clouds has interesting structures at three levels: 1) The “atomic” small-scale structure contains “crystals” whose faces are parallelograms or trapezoids, generalizing well-known examples such as (man:woman::king:queen), and get revealed when projecting out semantically irrelevant distractor features. 2) The “brain” intermediate-scale structure has significant spatial modularity; for example, math and code features form a “lobe” akin to functional lobes seen in neural fMRI images. 3) The “galaxy” scale large-scale structure of the feature point cloud is not isotropic, but instead has a power law of eigenvalues with steepest slope in middle layers. We hope that our findings serve as a stepping stone toward deeper understanding of SAE features and the workings of large language models.
@@ A.1 Co-occurrence measures
Definitions of co-occurrence based affinity measures: Let $n_{ij}$ be the number of times features $i$ and $j$ co-occur. Let $m_{11}$ be number of times $i$ and $j$ co-occur, $m_{00}$ be number of times $i$ and $j$ both do not occur, $m_{10}$ be number of times $i$ occurs but $j$ does not, $m_{1\bullet}$ be number of times $i$ occurs and $j$ either occurs or not, and so on. Then,
Jaccard similarity, Jaccard (1908):
@@
Dice score, Dice (1945):
@@
Simple matching coefficient:
@@
(a) Layer 0
(b) Layer 12
Figure 9: Histogram, over all features, of Phi coefficient with $k$-th nearest cosine similarity neighbor. Each line represents a different $k$. The “random” line is plotted by drawing a random feature for each feature, then computing the Phi coefficient. Features with higher cosine similarity have higher Phi coefficient, but this is less pronounced in layer 0 compared to layer 12.
@@ Appendix B Understanding Principal Components in Difference Space
Figure 11: Silhouette Score, a measure of clustering quality, as a function of reduced dimension in LDA. The plot indicates that training an affine transformation for semantic cluster separation is easier in mid-layer (layer 12), where the model starts to develop concept-level understanding of the input.
@@ Appendix C Breaking Down SAE Vectors by PCA Component
An additional investigation of structure we undertake is quantifying how SAE vectors are distributed throughout the PCA components of the activations vectors. To do this, we define a PCA score:
@@
Figure 12: Smoothed PCA scores for each SAE feature of the layer $12$, width $16k$, $L_{0}=176$ Gemma Scope 2b SAE, sorted by frequency. PCA score = $\frac{1}{n}\sum_{i}i*(pca_{i}@feature_{j})^{2}$, where $n$ is the number of PCA features. The smoothed curves just average this somewhat noisy metric over adjacent sorted features. This measures approximately where in the PCA each SAE feature lies, and shows that there is a dip into earlier PCA features in the last third of SAE features.

### v2

@@ 5.1 Shape Analysis
In RMT, the covariance matrix of $N$ random vectors from a multivariate Gaussian distribution follow a Wishart distribution Wishart (1928). Under this assumption, we would expect the eigenvalues to be relatively uniform or to follow the Marcenko–Pastur law Marchenko and Pastur (1967). In contrast, we observe a surprising derivation:
  • The eigenvalue spectrum of the point cloud decays as a power law rather than following the expected Wishart behavior.
  • As shown in Figure 6, this power law decay is more pronounced in SAE features compared to raw activations.
Since the abrupt drop off seen for the smallest eigenvalues is caused by limited data and vanishes in the limit $N\to\infty$, we dimensionally reduce the point cloud to its 100 largest principal components for all subsequent analysis in this section. We describe the shape of this high-dimensional point cloud as resembling a ”fractal cucumber”, whose width in successive dimensions falls off like a power law. We find such power law scaling is significantly less prominent for activations than for SAE features; it will be interesting for further work to investigate its origins.
Figure 7 (left) shows how the slope of the aforementioned power law depends on LLM layer, computed via linear regression against the 100 largest eigenvalues. We see a clear pattern where middle layers have the steepest power law slopes: (Layer 12 has slope --0.47, while early and late layers (e.g., Layers 0 and 24) have shallower slopes (--0.24 and --0.25), respectively. This may hint that middle layers act as a bottleneck, compressing information into fewer principal components, perhaps optimizing for more efficient representation of high-level abstractions. Figure 7 (right) compares the eigenvalue spectra of SAE features and neural activations, indicating a significantly steeper power law decay for SAE features. Activations, in contrast, exhibit a much slower decay, indicating weaker power law behavior and distinct geometric structures in the latent space. Figure 8 (left) explores the effective cloud volume (the determinant of the covariance matrix) of the point cloud, quantified by the log-determinant of the covariance matrix across layer. This volume variation further reflects the layer-specific changes in the structure and complexity of the latent space.
Figure 7: Eigenvalue distributions for SAE features and activations. Eigenvalues of the covariance matrix for SAE features (left) and neural activations (right) decay approximately as a power law, with slopes varying across layers. A scaled isotropic Gaussian spectrum is shown for comparison, highlighting the significantly steeper decay for SAE features. Eigenvalue spectra for activations show a much slower decay compared to SAE features, indicating weaker power law behavior and distinct geometric structures.
Figure 8: Layer-wise analysis of latent representations. (left): The power law slope ($\alpha$) of the eigenvalue spectrum (blue) and the log-determinant of the covariance matrix (green) vary across layers. Both metrics peak in intermediate layers, indicating significant structural changes in the latent space. (right): Estimated clustering entropy across layers with 95% confidence intervals. Middle layers exhibit reduced clustering entropy, while earlier and later layers show higher entropy, reflecting distributed and concentrated feature representations, respectively.
@@ 5.2 Clustering Analysis
Clustering of galaxies or microscopic particles is often quantified in terms of a power spectrum or correlation function. This is complicated for our very high-dimensional data, since the underlying density varies with radius and, for a high-dimensional Gaussian distribution, is strongly concentrated around a relatively thin spherical shell. For this reason, we instead quantify clustering by estimating the entropy of the distribution that the point cloud is assumed to be sampled from. We estimate the entropy $H$ from our SAE feature point cloud using the $k$-th nearest neighbor (k-NN) method Dasarathy (1991); Kozachenko and Leonenko (1987), computed as follows,
@@
where $r_{i}$ is the distance to the $k$-th nearest neighbor for point $i$, and $d$ is the dimensionality of the point cloud; $n$ is the number of points; the constant $\Psi$ is the digamma term from the k-NN estimation. As a baseline, the Gaussian entropy represents the maximum possible entropy for a given covariance matrix. For a Gaussian distribution with the same covariance matrix, the entropy is computed as follows:
@@
where $\lambda_{i}$ are the eigenvalues of the covariance matrix. We define the clustering entropy (often referred to as “negentropy” in physics as $H_{gauss}-H$, i.e., how much lower the entropy is than its maximum allowed value).
The estimated clustering entropy is shown in Figure 8 (right), plotted across different layers. The results indicate that the SAE point cloud is strongly clustered, particulary in the middle layers. This observation aligns with the reduced clustering entropy seen at intermediate layers, suggesting significant structural differences in the latent representations.
In future work, it will be interesting to investigate whether these variations depend mainly on the prominence of crystals or lobes in different layers, or have an altogether different origin (entirely different underlying mechanisms).
6 Conclusions
We have searched for structure in the SAE concept universe at three levels: (1) The “atomic” small-scale structure contains “crystals” whose faces are parallelograms or trapezoids, generalizing well-known examples such as (man:woman::king:queen), may be revealed when projecting out semantically irrelevant distractor features. (2) The “brain” intermediate-scale structure has significant spatial modularity; for example, math and code features form a “lobe” akin to functional lobes seen in neural fMRI images. (3) The “galaxy” large-scale structure of the feature point cloud is not isotropic, but instead has a power law of eigenvalues with steepest slope in middle layers.
While we have observed that SAE features exhibit geometric structure at multiple scales, we have not explained why this structure forms. We think that further work that not only studies the structure of SAE features, but also seeks to explain the origin of this structure, could be highly valuable. Such work may lead to refinements to our theory of how networks represent features in superposition or to insights that improve sparse autoencoder performance.
We hope that our findings serve as a stepping stone toward deeper understanding of SAE features and the workings of large language models, and that this deeper understanding will eventually help to improve the safety of AI systems as they continue to grow in power.
[…]
Conceptualization, J.E. and M.T.; software, E.J.M., D.D.B. and X.S.; formal analysis, Y.L., E.J.M. and D.D.B.; investigation, M.T.; writing—original draft preparation, Y.L., E.J.M., D.D.B., J.E., X.S. and M.T.; writing—review and editing, Y.L., E.J.M., D.D.B., J.E. and M.T.; visualization, Y.L., E.J.M. and D.D.B. All authors have read and agreed to the published version of the manuscript.
[…]
E.J.M., D.D.B, and M.T. are supported by IAIFI through NSF grant PHY-2019786. E.J.M. and J.E. are supported through the NSF GRFP (Grant No. 2141064). This work is supported by the Rothberg Family Fund for Cognitive Science.
[…]
Not applicable.
[…]
We provide code to replicate our results at this repository: https://github.com/ejmichaud/feature-geometry (accessed on 24 March 2025).
[…]
The authors declare no conflicts of interest.
[…]
@@ A.1 Co-Occurrence Measures
Definitions of co-occurrence-based affinity measures: Let $n_{ij}$ be the number of times features $i$ and $j$ co-occur. Let $m_{11}$ be number of times $i$ and $j$ co-occur, $m_{00}$ be number of times $i$ and $j$ both do not occur, $m_{10}$ be number of times $i$ occurs but $j$ does not, $m_{1\bullet}$ be number of times $i$ occurs and $j$ either occurs or not, and so on. Then, the following can be determined.
Jaccard similarity, Ref. Jaccard (1908), is as follows:
@@
Dice score, Ref. Dice (1945),is as follows:
@@
The simple matching coefficient is
@@
Figure 9: Histogram, over all features, of phi coefficient with $k$-th nearest cosine similarity neighbor for (left) Layer 0 and (right) Layer 12. Each line represents a different $k$. The “random” line is plotted by drawing a random feature for each feature, then computing the phi coefficient. Features with higher cosine similarity have higher phi coefficients, but this is less pronounced in Layer 0 compared to Layer 12.
@@ Appendix B Understanding Principal Components in Difference Space
Figure 11: Silhouette score, a measure of clustering quality, as a function of reduced dimension in LDA. The plot indicates that training an affine transformation for semantic cluster separation is easier in middle layers (Layer 12), where the model starts to develop concept-level understanding of the input.
@@ Appendix C Breaking Down SAE Vectors by PCA Component
An additional investigation of structure we undertake is quantifying how SAE vectors are distributed throughout the PCA components of the activations vectors. To accomplish this, we define a PCA score:
@@
Figure 12: Smoothed PCA scores for each SAE feature of Layer $12$, width 16k , $L_{0}=176$ Gemma Scope 2b SAE, sorted by frequency. PCA score = $\frac{1}{n}\sum_{i}i*(pca_{i}@feature_{j})^{2}$, where $n$ is the number of PCA features. The smoothed curves just average this somewhat noisy metric over adjacent sorted features. This measures approximately where in the PCA each SAE feature lies, and shows that there is a dip into earlier PCA features in the last third of SAE features.
{adjustwidth}
[…]